Project

General

Profile

Meeting 20-5-2015

Date: 120th may 2015
Time: 15:00 UK time
Attendance: Ray, Chis Hunt, Ed, Melissa, Chris Heidt
Apologies: Adam, Paul K,
h2. Agenda and Minutes

Actions from Last Meeting

  • Kalman doesn't work on straights. - Chris Hunt looking to fix. Kalman is in good shape. All basics in and small effects now being considered.
  • Real data unpacking broken. A proper summary of the issue, its status and timeline to be presented this week. - Paul.
  • Software update from Paul this meeting - PK. - Ongoing
  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick. - Ongoing
  • Add to John Govans' rota for daily walk round to test our setup. - Pierrick. - Ongoing
  • Look at docs to determine what to release as MICE notes and how - MU - ONGOING not highest priority
  • Update in depth Tracker schedule and MU Ed and Craig. - Happened schedule determined in documents 141, Schedule to Alan G by end of week.
  • Cryo lids still needed. Melissa to ask Geoff or Kevin. - temporarily covered Geoff to do permanent fix today and tomorrow - DONE
  • Put the above on the data section of the Tracker wiki. -MU. DONE doc 141
  • EO to speak to C. Heidt when ADC/VLPC and ADC calib and MAUS transition performing calibrations. - Scheduled for Friday

0. MDR3

The goals of MDR3 are (DAQ only):

1. Test the readout of the tracker with the DAQ pulser trigger

2. Integrate the tracker readout with the other PID detectors

3. Test the unpacking of the tracker data

Comments:
EO using online mon we will be able to check that we are getting data in the files.

1. Status of hardware, commissioning work and ongoing plan EO, AB, CMW, PR, MU

EO visit this week aimed to:
  • perform bias calibration scans(longer scan, more data).
  • check bias results.
  • Integrate remaining code with the tracker IOC.

2. Software: PK, AD, CH

In particular this week:
  • Data unpacking
  • Kalman
  • DAQ in data structure
  • CDB mapping and bias calibrations

Kalman: Chris Hunt. Kalman is in good shape For all tracks. All basics in and small effects now being considered. Multiple scattering is next weeks job. Residuals stored chisw and p vals being saved. Will recheck straight tracks.

Found recent bug which effects globals and is now fixed. with no large errors at start and end of run. In the trunk soon.

Geometry, Calibrations, Configurations, CDB - C. Heidt

C. Heidt: CDB met with Janusz script to up/download is done, needs to ad to the MAUS mapper to use the script as MUAS is run.

Monte Carlo

In a good state waiting on Durga for multiple tracks per Tracker.

Pattern Recognition - AD,

  • No method to get multiple tracks per trigger.

Online Displays - MU

  • 3/4 of the plots are now in and working and were used for mock data challenge. The other are the non-immediately-essentials and are coming soon.

Performance and Analysis -PK

Issues

Tracker Software Paper - AD, PK DR

  • C. Heidt: MC documentation coming along

Documentation - PK, AD, C. Heidt

C. Heidt: MC documentation coming along

4. Optimising the DAQ and Controls and Monitoring - EO, AB, PR

  • 50% done, but not enough time to fully go over before beam. Davids original code for MDR3 and new code to come in after first data.
  • CandM currently broken, being fixed next week. May be due to more complicated ioc. Not an issue for MDR3, but urgent.

5. MICE hall, inc Services hardware and requirements : CM

  • see schedule in documents.

6. Step IV Analysis Readiness - MU

  • Alignment data run requests submitted.
  • Unpacking and speed are biggest problems (PK)

7. Risks

7.1 Hardware

  • Daresbury electronics: vacuum, tracker interlocks
  • Reforming external waveguides (could lead to breakage)
  • Northside PRY (not us, but knock effect to us)

7.2 Software

  • Not yet able to test against multiple tracks per trigger with MC -Action for Durga
  • Real data unpacking broken DA

8. Tracker Documentation

We need to produce documentation for the control room and ISIS, for Steve B. We cannot necessarily get that far with operations and startup as we are not yet working with the CandM as we will but we can write something and finalise safety details.

8.1 Startup Procedure

8.2 Operations Procedure

8.3 Tracker Online Emergency Procedure

  • We have been asked to provide documentation for ISIS and eventually shifters in what to do in an emergency situation for the Tracker when running
    Documentation to Include:
    - schematics of the system
    - a list of associated checklists at power on, and who works through them.
    - a list of instructions to shifters on how to operate and monitor the system
    - emergency shutdown instructions
    - power-up instructions
    - trouble-shooting : what to watch for
    - a list of regular maintenance and safety checks

5. Other Items

  • Installation Manual - GB MU - ongoing
  • LED system - EO - External used for calib, Internal not used since Dec but will be used after gain and pedestals determined.
  • Hall probes - MU - Electronics readout devised and computing needs obtained. Cabling to be made.

AoB

New Actions

  • Kalman doesn't work on straights. - Chris Hunt looking to fix. Kalman is in good shape. All basics in and small effects now being considered.
  • Real data unpacking broken. A proper summary of the issue, its status and timeline to be presented this week. - Paul.
  • Software update from Paul this meeting - PK. - Ongoing
  • Prepare a tag out lock out procedure to stop people turning things off that we don't want them to (He etc. - Add the above to the controls and monitoring alarms so that if the above happens we know - Pierrick. - Ongoing
  • Add to John Govans' rota for daily walk round to test our setup. - Pierrick. - Ongoing
  • Look at docs to determine what to release as MICE notes and how - MU - ONGOING not highest priority
  • EO to speak to C. Heidt when ADC/VLPC and ADC calib and MAUS transition performing calibrations. - Scheduled for Friday
  • CDB script to up/download is done, needs to ad to the MAUS mapper to use the script as MUAS is run. - C. Heidt
  • Multiple tracks per Tracker. -Durga
  • Add online recon to trunk - MU
  • Fix C&M code - Pierrick
  • Check with DL to move interlocks after first user run - MU
  • Email update on Thursday of MDR3 result. - EO

Phone Details

  • The four digit company id: 4673 (callers from outside the UK)
  • The pin number: 419357

Phone numbers:

UK (Imperial College --preferred) 0845 545 4460 English
Austria 0820 4000 1503 English
0820 4000 15 02 German
Belgium 070 35 99 45 English
070 35 99 94 Dutch
070 35 98 66 French
Canada 1 712 432 2832 English
1 712 432 2844 French
Germany 01803 001 178 English
01570 248 8303 English (mobile)
01803 001 177 German
01570 248 8101 German (mobile)
Ireland 0818 270 007 English
Italy 848 390 166 English
848 391 819 Italian
Netherlands 0870 001 901 English
0870 001 909 Dutch
Poland 0801 003 543 English
0801 003 533 Polish
South Africa 087 550 0375 English
Spain 902 885 318 English
902 881 200 Spanish
Sweden 0939 2066 400 English
0939 2066 300 Swedish
Switzerland 0848 560 190 English
0848 560 378 French
0848 560 347 German
0848 560 385 Italian
United Kingdom (PowWowNow) 0844 4 73 73 73 English
United States 1 415 363 0833 English